Calculate the number of oxygen molecules required to convert 7.50dm^3 of SO2 measured at s.t.p to SO4
2so4[g]+o2[g]-2so3[g[
(Molar volume of gas=22.4dm^3,Avogadro's number=6.02×10^23)
1
Expert's answer
2020-01-08T02:02:14-0500

7.5dm^3=7.5L

n(SO2)=7.5L/22.4L=0.3348mol

n(O2)=0.3348mol/2=0,1674mol

N(O2)=0,1674*6.02×10^23=1.007*10^23